# AV-RIR: Audio-Visual Room Impulse Response Estimation

### v1

@@ 4.3 Perceptual Evaluation
We perform a perceptual evaluation to evaluate how well our estimated RIR captures target environment acoustics (Table 3). For each environment, we provide the ground truth (GT) speech, generated speech using Image2Reverb [81], VAM [13], our AV-RIR and the environment image. We asked the participants to select the generated speech that sounds closer to the GT speech.
Initial Pre-Screening : As the initial pre-screening, the questions were designed to test whether the participants had normal hearing and proper listening devices. In our first question (Q1), we replaced one of the generated speeches with a GT speech to see whether the participant chose the GT speech. We counted the results of the participants who correctly chose GT speech among generated speech in Q1.
We selected 6 scenes with varying complexity with $T_{60}$ ranging from 0.2 seconds to 0.7 seconds. We can see that, irrespective of the $T_{60}$ and the environment complexity, 56% to 79% of the participants said that the generated speech from AV-RIR closely matched the GT speech.
@@ 5 Conclusion, Limitations and Future Work
AV-RIR assumes stationary single-talker input speech or single-source audio without noise. Future work aims to tackle multi-channel RIR estimation and RIR estimation from noisy, multi-source environments with moving sources.
6 Supplementary:
@@
  • The qualitative results of our AV-RIR via a supplementary video.
@@ 6.1 Supplementary Video
We provide a supplementary video  [^2] https://www.youtube.com/watch?v=tTsKhviukAE showing the qualitative results of RIR estimation with AV-RIR when applied to three different tasks. In addition, we compare the enhanced speech from our AV-RIR with ground truth clean speech. We also demonstrate our approach’s failure cases in the supplementary video.
@@ 6.2 Far-field Automatic Speech Recognition
Training Dataset Word Error Rate normal-$\downarrow$ [%]

### v2

@@ 4.3 Perceptual Evaluation
In Table 3 we report scores for perceptual evaluation of our estimated RIRs. For each environment, we provide the ground truth (GT) speech, generated speech using Image2Reverb [81], VAM [13], our AV-RIR and the environment image. The participants were asked to select the generated speech that sounds closer to the GT speech.
Deatils on initial pre-screening of participants is described in the Appendix. We select 6 scenes with varying complexity with $T_{60}$ ranging from 0.2 seconds to 0.7 seconds. We can see that, irrespective of the $T_{60}$ and the environment complexity, 56% to 79% of the participants said that the generated speech from AV-RIR closely matched the GT speech.
@@ 5 Conclusion, Limitations and Future Work
AV-RIR assumes stationary single-talker input speech or single-source audio without noise. Future work aims to tackle multi-channel RIR estimation and RIR estimation from noisy, multi-source environment with moving sources.
6 Table of Contents:
@@
  • The qualitative results of our AV-RIR via a supplementary video [^2] https://www.youtube.com/watch?v=tTsKhviukAE.
@@ 6.1 Supplementary Video
We provide a supplementary video showing the qualitative results of RIR estimation with AV-RIR when applied to three different tasks. In addition, we compare the enhanced speech from our AV-RIR with ground truth clean speech. We also demonstrate our approach’s failure cases in the supplementary video.
@@ 6.2 Far-field Automatic Speech Recognition
Training Dataset Word Error Rate $\downarrow$ [%]